OpenStudy (anonymous):

3 children are camping out & go to sleep with a pack of marshmallow unopened in their tent. ONe child wakes up & eats 1/4 of the marshmallows & then goes back to sleep. A short time later another child wakes up & eats 1/3 of the remaining marshmallows & then goes back to sleep. Later in the night the other child wakes up & eats 1/2 of the remaining marshmallows & then goes back to sleep. When they all awake in the morning there is only 9 marshmallow left. How many marshmallows were in the pack at the start of the night? What is the simplest method for me to work this problem out?

OpenStudy (mathstudent55):

Original number of marshmallows: m First child eats 1/4: eats (1/4)m, and (3/4)m are left Second child eats (1/3)(3/4)m, (3/4)m - (1/3)(3/4)m are left = (1/2)m Third child eats (1/2)(1/2)m, (1/2)m - (1/4)m are left = (1/4)m (1/4)m = 9 m = 36

OpenStudy (anonymous):



OpenStudy (anonymous):

ok, thank you

OpenStudy (mathstudent55):

wlcm